OpenStudy (anonymous):

value of x where the tangent line to the graph of f(x)=-x^2+8x+2 is horizontal.

OpenStudy (anonymous):

i need exp.

OpenStudy (anonymous):

Take the derivative and set it to zero. Use: \[\frac{d}{dx}x^n = nx^{n-1}\]

OpenStudy (anonymous):

2x-8? like that

OpenStudy (anonymous):

It would be -2x+8=0

OpenStudy (anonymous):

+

OpenStudy (anonymous):

then ill divide it?

OpenStudy (anonymous):

Yeah, so: -2x=-8 x=4

OpenStudy (anonymous):

oh i get it! ;) thanks soo much.